OpenStudy (anonymous):

-3x Is less than or equal to 27 A. x -9 B. x -9 C. x > 9 D. x = -9 E. x 9 F. x < -9

OpenStudy (anonymous):

d

OpenStudy (anonymous):

\[A. x \ge -9 \] \[B. X \le -9\] \[E. X \le 9\]

OpenStudy (anonymous):

A

jhonyy9 (jhonyy9):

B is sure right

OpenStudy (anonymous):

It was A Thank you both for helping!!